Accelerator Update for March 7, 2008 - March 10, 2008
Friday March 7
The day shift began with Operations monitoring the Tevatron (TeV) store 5965 (32.08E30) and the Recycler stash (276.1E10), with the Antiproton Source stacking (43.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, with NuMI off for horn repair, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
Operations terminated store 5965 at 8:34 AM.
Operations put the TeV into shot setup at 8:57 AM.
At 9:20 AM, a Linac RF station (LRF1) tripped off. It reset.
Operations established store 5966 at 10:42 AM with an initial luminosity of 239.28E30.
The Neutron Therapy Facility (NTF) reported at 1:04 PM that they had completed treatment on their last patient for the day.
Operations put Pbar into standby at 1:29 PM in preparation for an access to check on the lithium lens cooling system. They came out of access about an hour later.
A MI expert began a study at 3:18 PM.
Operations reported at 3:56 PM that the Recycler had stashed antiprotons twice during the day shift.
The MI study ended at 4:49 PM.
At 7:12 PM, a Booster vacuum pump (IP111) tripped off. It reset.
At 8:48 PM, a Linac Klystron RF station (KRF4) tripped off. It reset.
At 8:57 PM, Operations reported finding five vacuum pumps off in the MI-8 transfer line. Operators turned the pumps on, but then a vacuum burst caused the pumps to trip off again. Operators reset everything, but left two of the pumps off (IP831 and IP832).
At 10:38 PM, KRF3 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Saturday March 8
The midnight shift began with Operations monitoring store 5966 (63.27E30) and the stash (222.3E10), with Pbar stacking (29.6E10), with NuMI off for horn repair, and with BNB taking beam.
At 2:51 AM, three MI RF vacuum pumps (604E, 605A & B) tripped off. They all reset.
Operations terminated store 5966 at 4:04 AM.
Operations prepared D0 and CDF for access at 4:15 AM.
At 5:43 AM, MIRF5 tripped off. It reset.
Operations reported at 8 AM that the Recycler had stashed four times during the midnight shift.
D0 came out of access at 8:13 AM.
CDF came out of access at 9:10 AM.
A NuMI expert asked for a search and secure of MI-65 at 9:34 AM.
Operations reported at 11:20 AM that they had completed their search and secure and had a NuMI beam permit.
At 11:25 AM, the AP1 transfer line tripped off due to bad beam positions. It reset.
Operations put the TeV through a dry squeeze at 11:36 AM.
A power supply expert reported at 12:40 PM that the NuMI quadrupole power supplies 101 & 106 had some regulator fuse problems. He brought them back online.
Operations reported at 12:42 PM that the TeV needed an orbit smooth. They called in a TeV expert.
At 12:43 PM, Operations reported that the AP1 line was continuing to trip off due to a problem with over-temperature trips on the HV100 magnets. Operations called in an expert.
Power supply experts reported at 1:38 PM that they had adjusted the NuMI quadrupole supply’s (E:QD111) ground fault sensitivity.
At 2:26 PM, a Linac quadrupole bulk power supply (Q6) began tripping off on a temperature indication. It would not stay reset. Operations called in a Linac expert.
Operations put the TeV into shot setup at 3:31 PM.
A Linac expert reported at 5:42 PM that he had Q6 bulk power supply back online. Operations resumed sending beam to NuMI and BNB.
Operations established store 5968 at 7:23 PM with an initial luminosity of 236.04E30.
A MI expert began tuning for seven-turn running for NuMI at 7:48 PM.
Operations reported at 8 PM that the Recycler had not stashed during the last 12 hours.
A Pbar RF expert reported at 9:28 PM that ARF2 had a bad driver and he had fixed it.
At 10:55 PM, a driver who was filling a MuCool liquid nitrogen tank reported that the tank had a leak due to a bad valve. The Cryo Coordinator said cryo system technicians would empty this tank and repair the valve.
Sunday March 9
The midnight shift began with Operations monitoring store 5968 and the stash, with Pbar stacking, and with NuMI and BNB taking beam.
At 1:25 AM, KRF2 tripped off. It reset.
At 3:03 AM, BRF11 tripped off. It reset.
At 6:02 AM, KRF2 tripped off. It reset.
At 6:41 AM, Operations reported that a Booster quadrupole power supply (B:QL1) was not tracking properly.
Operations reported at 7:57 AM that the Recycler had stashed six times during the last 12 hours.
At 8:33 AM, a Booster correction magnet power supply (VL1) had tripped its breaker. It would reset, but then immediately trip back off. Operations contacted experts.
At 8:54 AM, KRF3 tripped off. It reset.
Operations prepared Booster for an access at 11:15 AM.
Booster came out of access at 12:35 PM.
Booster experts reported at 12:42 PM that they found a loose power lead for VL1. The experts also changed out regulators and control voltage modules for a number of other power supplies.
Booster resumed sending beam at 12:46 PM.
At 12:54 PM, MP02 tripped off for the first of four times in 16 minutes. Operations contacted an expert.
At 2:12 PM, KRF6 tripped off. It reset.
Operations terminated store 5968 at 2:46 PM.
Operations prepared CDF for an access at 3:02 PM.
Operations reported at 7:57 PM that the Recycler had stashed twice during the last 12 hours.
CDF came out of access at 8:27 PM.
Operations put the TeV through a dry squeeze at 8:48 PM.
Operations put the TeV in shot setup at 9:12 PM.
Operators completed a Booster injection tune at 9:24 PM
About 9:57 PM, a BNB expert came in to access and make up their horn LCW system. This took about 30 minutes.
BNB resumed taking beam at 10:53 PM.
Operations established store 5970 at 11:01 PM with an initial luminosity of 263.9E30.
At 11:08 PM, a Klystron vernier tripped off on a transformer LCW fault for the first of six times in 23 minutes.
Monday March 10
The midnight shift began with Operations monitoring store 5970 and the stash, with Pbar stacking, and with NuMI and BNB taking beam.
At 1:37 AM, BRF14 tripped off. It reset.
At 4:28 AM, KRF3 tripped off. It reset.
Operations turned off the NuMI beam at 5:05 AM for a scheduled shutdown to work on their dehumidifiers.
At 5:56 AM, the MI-8 transfer line suffered a vacuum burst. The vacuum had mostly recovered by 6:20 AM.
Operations had NuMI prepared for access by 7:45 AM.
Operations reported at 8 AM that the Recycler had stashed five times during the last 12 hours.
The Plans for Monday
The plans for today are to stack, stash, and store. NuMI will resume taking beam on Tuesday evening (3/11/08).
Comments and Suggestions
|last modified 03/10/2008 email Fermilab|
|Security, Privacy, Legal|